40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Dec 4 21:48:01 2009, Jenne, Update, oplevs, Oplevs centered, IP_POS and IP_ANG centered Oplev_IPang_screenshot_4Dec2009.png
    Reply  Fri Dec 4 23:17:55 2009, rob, Update, oplevs, Oplevs centered, IP_POS and IP_ANG centered 
       Reply  Sat Dec 5 01:40:11 2009, Koji, Update, oplevs, Oplevs centered, IP_POS and IP_ANG centered 
          Reply  Sat Dec 5 18:23:48 2009, Koji, Update, oplevs, Oplevs centered, IP_POS and IP_ANG centered Screenshot_091205_1830.png
Message ID: 2358     Entry time: Sat Dec 5 18:23:48 2009     In reply to: 2354
Author: Koji 
Type: Update 
Category: oplevs 
Subject: Oplevs centered, IP_POS and IP_ANG centered 

NOTE: HEPA is on at its full.

[[[OK]]] Align the suspended optics (by Rob)
[[[OK]]]
Align the oplevs again
[[[OK]]] Take snapshots for the suspensions/QPDs/IO QPDs/PZT strain gauges
[[[OK]]] Align the IP_POS, IP_ANG
[[[OK]]] Align the PSL table QPDs, the MC WFS QPDs, and the MCT QPD
[[[OK]]] Align the aux laser for the absolute length 


Alignment of the aux laser

o Go to only ITMX mode:
Save the alignment of the mirrors. Activate X-arm mode. Misalign ITMY and ETMX.

o Inject the aux beam:
Open the shutter of the aux NPRO. Turn the injection flipper on.

o Look at the faraday output:
There are several spots but only one was the right one. Confirm the alignment to the thorlabs PD. Connect the oscilloscope to the PD out with a 50Ohm termination.
Thanks to the Alberto's adjustment, the beat was already there at around 10MHz. After the PD adjustment, the DC was about 600mV, the beat amplitude was about 50mVpp.

o Adjust the aux beam alignment:
Adjust the alignment of the aux beam by the steering mirrors before the farady isolator. These only change the alignment of the aux beam independently from the IFO beam.
After the alignment, the beat amplitude of 100mVpp was obtained.

o Closing
Close the shutter of the NPRO. Turn off the flipper mirror. Restore the full alignment of the IFO.

Attachment 1: Screenshot_091205_1830.png  288 kB  | Hide | Hide all
Screenshot_091205_1830.png
ELOG V3.1.3-